# Gaussian-Process-based Adaptive Tracking Control with Dynamic Active Learning for Autonomous Ground Vehicles

### v1

@@ Abstract
This article proposes an active-learning-based adaptive trajectory tracking control method for autonomous ground vehicles to compensate for modeling errors and unmodeled dynamics. The nominal vehicle model is decoupled into lateral and longitudinal subsystems, which are augmented with online Gaussian Processes (GPs), using measurement data. The estimated mean functions of the GPs are used to construct a feedback compensator, which, together with an LPV state feedback controller designed for the nominal system, gives the adaptive control structure. To assist exploration of the dynamics, the paper proposes a new, dynamic active learning method to collect the most informative samples to accelerate the training process. To analyze the performance of the overall learning tool-chain provided controller, a novel iterative, counterexample-based algorithm is proposed for calculating the induced $\mathcal{L}_{2}$ gain between the reference trajectory and the tracking error. The analysis can be executed for a set of possible realizations of the to-be-controlled system, giving robust performance certificate of the learning method under variation of the vehicle dynamics. The efficiency of the proposed control approach is shown on a high-fidelity physics simulator and in real experiments using a 1/10 scale F1TENTH electric car.
@@ I Introduction
The state-of-the-art trajectory tracking algorithms for ground vehicles are usually studied in the context of autonomous racing [1], utilizing both model-free [2, 3] and model-based [4, 5] approaches. While model-free methods have gained significant attention, in the context of autonomous vehicles, model-based solutions are still often favored due to safety concerns. The performance of model-based approaches is dependent on the accuracy of the underlying model, therefore, precise identification of the motion dynamics is required. For this purpose, often first-principles-based models are used, but even after estimating the parameters of such models, unknown and hard-to-model dynamic effects that are specific to each vehicle often have significant effects on the behavior of the system. As a result, model uncertainties are inevitable and can lead to performance degradation in precise maneuvering in practice. To efficiently mitigate the effect of modeling uncertainties, adaptive algorithms have been extensively studied, which either estimate the parameters of the system dynamics, [6], or directly adapt the parameters of a control policy [7]. Furthermore, model augmentation, where a nominal vehicle model, typically based on first principles, is complemented with a learning component to capture the residual model mismatch, has been investigated. These methods usually utilize machine learning in terms of artificial neural networks (ANN) [8], or Gaussian processes (GP) [9] to capture the unknown model dynamics. GP augmentation has proven to be beneficial for a wide range of mobile robotic applications such as cars [10] quadcopters [11] and robotic arms [12], due to its high approximation capability and the uncertainty characterization of the estimates.
@@
For GP-augmented models, nonlinear model predictive control (NMPC) algorithms are usually preferred [9, 10, 11, 12]. However, due to the online optimization, real-time NMPCs require substantial computational resources compared to classical feedback methods, which are still applied in areas, where processing power is limited [5]. Therefore, this article proposes an effective GP-based compensation method with low computational complexity to adapt to modeling uncertainties.
@@
Collecting measurement data for the training of the GPs in real-world scenarios is crucial for the effectiveness of the learning methods and it can be a time and energy-consuming task. Often, this step is overlooked in the literature. Furthermore, constraints of the specific environments (e.g. limited test areas, obstacles) and the motion dynamics of the vehicle make the data acquisition more challenging. Active learning methods have been a popular choice for the identification of unknown dynamics [13] and to refine GP models [14, 15] using their posterior distribution. However, existing methods assume that input-output training data can be easily collected, i.e., the function to be approximated can be evaluated at any value of its arguments. In our case, the inputs of the GP are the states of the motion dynamics. Therefore, to collect training data, the dynamic model has to be driven to the required state configuration. To address this issue, we propose a novel dynamic active learning approach that takes into account the system dynamics and systematically designs experiments for data collection.
Lastly, providing performance guarantees is essential in autonomous robotic applications. Feedback algorithms for autonomous cars usually rely on decoupled and simplified dynamic models [16, 17] or are based on only kinematics [18], therefore, stability for the nonlinear closed-loop can not be guaranteed by design. However, computing stability guarantees for nonlinear systems is difficult, especially with learning components such as GPs or ANNs in the closed loop.
@@ II-A Gaussian Processes
To estimate a scalar unknown function $f:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}$ with a GP, let $X^{N}=[x_{1}^{\top}\;\cdots\;x_{N}^{\top}]\in\mathbb{R}^{N\times n_{x}}$ denote the collection of $N$ number of input values and let $Y^{N}=[y_{1}\;\cdots\;y_{N}]^{\top}\in\mathbb{R}^{N}$ be the corresponding noisy output measurements, forming the data set $\mathcal{D}^{N}=\{X^{N},Y^{N}\}$, generated by the following model:
$y_{i}=f(x_{i})+\epsilon_{i},\quad i\in\mathbb{I}_{1}^{N},$ (1)
@@
In terms of definition, a scalar-valued Gaussian Process $\mathcal{GP}$ assigns to every point $x\in\mathbb{R}^{n_{x}}$ a random variable $\mathcal{GP}(x)$, such that, for any finite set $\{x_{i}\}_{i=1}^{N}\subset\mathbb{R}^{n_{x}}$, the joint probability distribution of $\mathcal{GP}(x_{1}),\cdots,\mathcal{GP}(x_{N})$ is Gaussian. Due to this property, $g\sim\mathcal{GP}(m,\kappa)$ is fully determined by its mean $m$ and covariance $\kappa$ expressed as
$m(x)$ $=\mathbb{E}[g(x)],$ (2a)
$\kappa(x,\tilde{x})$ $=\mathbb{E}[(g(x)-m(x))(g(\tilde{x})-m(\tilde{x}))],$ (2b)
where $x,\tilde{x}\in\mathbb{R}^{n_{x}}$ and $\mathbb{E}$ is the expectation. This distribution describes our prior belief of the function space in which estimate of the the unknown function $f$ is searched for. In the remainder of the paper, we assume w.l.o.g. that the prior mean is zero ($m(x)=0$) and covariance of the distribution is characterized with a squared exponential (SE) kernel, a common choice for estimation of smooth functions:
$\kappa_{\mathrm{SE}}(x,\tilde{x})=\sigma_{\mathrm{f}}^{2}\exp\left(-\frac{1}{2}(x-\tilde{x})^{\top}\Lambda^{-1}(x-\tilde{x})\right),$ (3)
where $\sigma_{\mathrm{f}}\in\mathbb{R}$ is a scaling factor and $\Lambda\in\mathbb{R}^{n_{x}\times n_{x}}$ is a positive definite and symmetric matrix that determines the smoothness of the candidate functions. The parameters of the considered kernel function, i.e., $\sigma_{\mathrm{f}}$, $\Lambda$ and $\sigma_{\epsilon}$ in (1) are the so-called hyperparameters of the prior distribution and they are collected in the vector $\theta\in\mathbb{R}^{n_{\theta}}$. Note that other kernels can also be considered as it is discussed in [21].
@@
is the probability density function of the $Y^{N}$ outputs seen as random variables conditioned on $X^{N}$ and $\theta$, where $[K_{NN}]_{i,j}=\kappa(x_{i},x_{j}),\;i,j\in\mathbb{I}_{1}^{N}$. Simple derivation leads to that the predictive distribution for $g(x_{\star})$ at test point $x_{\star}$ is the posterior $p(g(x_{\star})|\mathcal{D}^{N},x_{\star})=\mathcal{N}(\mu(x_{\star}),\Sigma(x_{\star}))$, with
$\mu(x_{\star})$ $=K_{N}^{\top}(x_{\star})(K_{NN}+\sigma_{\epsilon}^{2}I)^{-1}Y^{N},$ (5a)
$\Sigma(x_{\star})$ $=\kappa(x_{\star},x_{\star})-K_{N}^{\top}(x_{\star})(K_{NN}+\sigma_{\epsilon}^{2}I)^{-1}K_{N}(x_{\star}),$ (5b)
where $[K_{N}(x_{\star})]_{i}=\kappa(x_{i},x_{\star}),\;i\in\mathbb{I}_{1}^{N}$. Eq. (5a) characterizes the mean as the approximation of the unknown function $f$ and (5b) is the variance, which gives the uncertainty of the approximation.
@@ II-B Sparse Gaussian Processes
One way to tackle the computational limitation of GP regression is to use sparse GPs (SGPs), where the number of data points used in the evaluation is limited to a fixed number. The goal of the SGP is to create a virtual dataset (often referred to as inducing points) defined as $\mathcal{D}^{M}=\{X^{M},Y^{M}\}$, where $X^{M}=[\hat{x}_{1}^{\top}\cdots\hat{x}_{M}^{\top}]\in\mathbb{R}^{M\times n_{x}}$ and $Y^{M}=[\hat{y}_{1}\cdots\hat{y}_{M}]\in\mathbb{R}^{M}$ such that the Kullback-Leiber (KL) divergence between the posteriori distribution obtained from $\mathcal{D}^{M}$ and the original distribution based on $\mathcal{D}^{N}$ is minimal, while $M\ll N$.
@@
$\mathcal{L}(\mathcal{D}^{N},X^{M},\theta)=\frac{N}{2}\log(2\pi)+\frac{1}{2}Y^{\top}(W_{NN}+\sigma_{\epsilon}^{2}I_{N})Y\\ +\frac{1}{2}\log\det(W_{NN}+\sigma_{\epsilon}^{2}I_{N})+\frac{1}{2}\mathrm{tr}(K_{NN}-W_{NN})$ (8)
where $W_{NN}=K_{NM}K_{MM}^{-1}K_{MN}$ with $[K_{MN}]_{i,j}=\kappa({\hat{x}_{i}},x_{j}),\;i\in\mathbb{I}_{1}^{M},\;j\in\mathbb{I}_{1}^{N}$, which is the covariance matrix between the pseudo inputs and all training inputs, $[K_{MM}]_{i,j},\;i,j=\mathbb{I}_{1}^{M}$ is the covariance of the pseudo inputs. By maximizing (8), the hyperparameters of the GP and the pseudo inputs can be jointly obtained, i.e.
@@
After the training, the resulting posterior distribution at an arbitrary test point $x_{\star}$ is $\mathcal{N}(\mu_{\mathrm{p}}(x_{\star}),\Sigma_{\mathrm{p}}(x_{\star}))$, with
$\mu_{\mathrm{p}}(x_{\star})$ $=K_{M}^{\top}(x_{\star})Q_{MM}K_{MN}\sigma^{-2}_{\epsilon}Y^{N},$ (10a)
$\begin{split}\Sigma_{\mathrm{p}}(x_{\star})&=\kappa(x_{\star},x_{\star})-W_{MM}(x^{\star})+K_{M}^{\top}(x^{\star})K_{MM}^{-1}K_{M}(x^{\star}),\end{split}$ (10b)
where $[K_{M}(x_{\star})]_{i}=\kappa(x_{i},x_{\star}),\;i\in\mathbb{I}_{1}^{M}$ and $W_{MM}(x_{\star})=K_{M}^{\top}(x\star)K_{MM}^{-1}K_{M}(x_{\star})$, which corresponds to a Nyström projection of the original GP to the pseudo inputs $X^{M}$. This formulation reduces the $\mathcal{O}(N^{3})$ computational complexity of the evaluation of the mean to $\mathcal{O}(M)$ and the training to $\mathcal{O}(NM^{2})$. In the following sections, SGPs will be utilized as the learning component of the proposed adaptive control approach.

### v2

@@ Abstract
This article proposes an active-learning-based adaptive trajectory tracking control method for autonomous ground vehicles to compensate for modeling errors and unmodeled dynamics. The nominal vehicle model is decoupled into lateral and longitudinal subsystems, which are augmented with online Gaussian Processes (GPs) using measurement data. The estimated mean functions of the GPs are used to construct a feedback compensator, which, together with an LPV state feedback controller designed for the nominal system, gives the adaptive control structure. To assist exploration of the dynamics, the paper proposes a new, dynamic active learning method to collect the most informative samples to accelerate the training process. To analyze the performance of the overall learning tool-chain provided controller, a novel iterative, counterexample-based algorithm is proposed for calculating the induced $\mathcal{L}_{2}$ gain between the reference trajectory and the tracking error. The analysis can be executed for a set of possible realizations of the to-be-controlled system, giving robust performance certificate of the learning method under variation of the vehicle dynamics. The efficiency of the proposed control approach is shown on a high-fidelity physics simulator and in real experiments using a 1/10 scale F1TENTH electric car.
@@ I Introduction
The state-of-the-art trajectory tracking algorithms for ground vehicles are usually studied in the context of autonomous racing [1], utilizing both model-free [2] and model-based [3, 4, 5] approaches. While model-free methods have gained significant attention, in the context of autonomous vehicles, model-based solutions are still often favored due to safety concerns. The performance of model-based approaches is dependent on the accuracy of the underlying model, therefore, precise identification of the motion dynamics is required. For this purpose, often first-principles-based models are used, but even after estimating the parameters of such models, unknown and hard-to-model dynamic effects that are specific to each vehicle often have significant effects on the behavior of the system. As a result, model uncertainties are inevitable and can lead to performance degradation in precise maneuvering in practice. To efficiently mitigate the effect of modeling uncertainties, adaptive algorithms have been extensively studied, which either estimate the parameters of the system dynamics, [6], or directly adapt the parameters of a control policy [7]. Furthermore, model augmentation, where a nominal vehicle model, typically based on first principles, is complemented with a learning component to capture the residual model mismatch, has been investigated. These methods usually utilize machine learning in terms of artificial neural networks (ANN) [8], or Gaussian processes (GP) [9] to capture the unknown model dynamics. GP augmentation has proven to be beneficial for a wide range of mobile robotic applications such as cars [10] quadcopters [11] and robotic arms [12], due to its high approximation capability and the uncertainty characterization of the estimates.
@@
For GP-augmented models, nonlinear model predictive control (NMPC) algorithms are usually preferred [9, 10, 11, 12]. However, due to the online optimization, real-time NMPCs require substantial computational resources compared to classical feedback methods, therefore, the latter is still favored in applications, where processing power is limited [4]. Therefore, this article proposes an effective GP-based feedback compensation method with low computational complexity to adapt to modeling uncertainties.
@@
Collecting measurement data for the training of the GPs in real-world scenarios is crucial for the effectiveness of the learning methods and it can be a time and energy-consuming task. Often, this step is overlooked in the literature. Furthermore, constraints of the specific environments (e.g. limited test areas, obstacles) and the motion dynamics of the vehicle make the data acquisition challenging. Active learning methods have been a popular choice for the identification of unknown dynamics [13] and to refine GP models [14, 15] using their posterior distribution. However, existing methods assume that input-output training data can be easily collected, i.e., the function to be approximated can be evaluated at any value of its arguments. In our case, the inputs of the GP are the states of the motion dynamics. Therefore, to collect training data, the dynamic model has to be driven to the required state configuration. To address this issue, we propose a novel dynamic active learning approach that takes into account the system dynamics and systematically designs experiments for data collection.
Providing performance guarantees is a crucial requirement in the study of autonomous systems. In practice, feedback algorithms for autonomous vehicles often rely on decoupled or simplified dynamic models [16, 17], or are based solely on kinematic representations [18]. As a result, these approaches cannot ensure stability of the full nonlinear closed-loop system by design. Guaranteeing stability for nonlinear systems is inherently challenging, particularly when learning-based components, such as GPs or ANNs, are included in the control loop [19].
@@ II-A Gaussian Processes
To estimate a scalar unknown function $f:\mathbb{R}^{n_{\mathrm{x}}}\rightarrow\mathbb{R}$ with a GP, let $X^{N}=[{{\color[rgb]{0,0,0}\mathpzc{x}}}_{1}^{\top}\;\cdots\;{{\color[rgb]{0,0,0}\mathpzc{x}}}_{N}^{\top}]\in\mathbb{R}^{N\times n_{\mathrm{x}}}$ denote the collection of $N$ number of input values and let $Y^{N}=[{{\color[rgb]{0,0,0}\mathpzc{y}}}_{1}\;\cdots\;{{\color[rgb]{0,0,0}\mathpzc{y}}}_{N}]^{\top}\in\mathbb{R}^{N}$ be the corresponding noisy output measurements, forming the data set $\mathcal{D}^{N}=\{X^{N},Y^{N}\}$, generated by the following model:
${{\color[rgb]{0,0,0}\mathpzc{y}}}_{i}=f({{\color[rgb]{0,0,0}\mathpzc{x}}}_{i})+\epsilon_{i},\quad i\in\mathbb{I}_{1}^{N},$ (1)
@@
In terms of definition, a scalar-valued Gaussian Process $\mathcal{GP}$ assigns to every point ${{\color[rgb]{0,0,0}\mathpzc{x}}}\in\mathbb{R}^{n_{\mathrm{x}}}$ a random variable $\mathcal{GP}({{\color[rgb]{0,0,0}\mathpzc{x}}})$, such that, for any finite set $\{{{\color[rgb]{0,0,0}\mathpzc{x}}}_{i}\}_{i=1}^{N}\subset\mathbb{R}^{n_{\mathrm{x}}}$, the joint probability distribution of $\mathcal{GP}({{\color[rgb]{0,0,0}\mathpzc{x}}}_{1}),\cdots,\mathcal{GP}({{\color[rgb]{0,0,0}\mathpzc{x}}}_{N})$ is Gaussian. Due to this property, $g\sim\mathcal{GP}(m,\kappa)$ is fully determined by its mean $m$ and covariance $\kappa$ expressed as
$m({{\color[rgb]{0,0,0}\mathpzc{x}}})$ $=\mathbb{E}[g({{\color[rgb]{0,0,0}\mathpzc{x}}})],$ (2a)
[…]
where ${{\color[rgb]{0,0,0}\mathpzc{x}}},\tilde{{{\color[rgb]{0,0,0}\mathpzc{x}}}}\in\mathbb{R}^{n_{\mathrm{x}}}$ and $\mathbb{E}$ is the expectation. This distribution describes our prior belief of the function space in which the estimate of the the unknown function $f$ is searched for. In the remainder of the paper, we assume w.l.o.g. that the prior mean is zero ($m({{\color[rgb]{0,0,0}\mathpzc{x}}})=0$) and covariance of the distribution is characterized with a squared exponential (SE) kernel, a common choice for estimation of smooth functions:
[…]
where $\sigma_{\mathrm{f}}\in\mathbb{R}$ is a scaling factor and $\Lambda\in\mathbb{R}^{n_{\mathrm{x}}\times n_{\mathrm{x}}}$ is a positive definite and symmetric matrix that determines the smoothness of the candidate functions. The parameters of the considered kernel function, i.e., $\sigma_{\mathrm{f}}$, $\Lambda$ and $\sigma_{\epsilon}$ in (1) are the so-called hyperparameters of the prior distribution and they are collected in the vector $\theta\in\mathbb{R}^{n_{\theta}}$. Note that other kernels can also be considered, as it is discussed in [22].
@@
is the probability density function of the $Y^{N}$ outputs seen as random variables conditioned on $X^{N}$ and $\theta$, where $K_{NN}]_{i,j}=\kappa({{\color[rgb]{0,0,0}\mathpzc{x}}}_{i},{{\color[rgb]{0,0,0}\mathpzc{x}}}_{j}),\;i,j\in\mathbb{I}_{1}^{N}$. Simple derivation leads to that the predictive distribution for $g({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star})$ at test point $\color[rgb]{0,0,0}\mathpzc{x}$$_{\star}$ is the posterior $p(g({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star})|\mathcal{D}^{N},{{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star})=\mathcal{N}(\mu({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star}),\Sigma({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star}))$, with
$\mu({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star})$ $=K_{N}^{\top}({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star})(K_{NN}+\sigma_{\epsilon}^{2}I)^{-1}Y^{N},$ (5a)
$\Sigma({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star})$ $=\kappa({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star},{{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star})-K_{N}^{\top}({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star})(K_{NN}+\sigma_{\epsilon}^{2}I)^{-1}K_{N}({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star}),$ (5b)
where $K_{N}({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star})]_{i}=\kappa({{\color[rgb]{0,0,0}\mathpzc{x}}}_{i},{{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star}),\;i\in\mathbb{I}_{1}^{N}$. Eq. (5a) characterizes the mean as the approximation of the unknown function $f$ and (5b) is the variance, which gives the uncertainty of the approximation.
@@ II-B Sparse Gaussian Processes
One way to tackle the computational limitation of GP regression is to use sparse GPs (SGPs), where the number of data points used in the evaluation is limited to a fixed number. The goal of the SGP is to create a virtual dataset (often referred to as inducing points) defined as $\mathcal{D}^{M}=\{X^{M},Y^{M}\}$, where $X^{M}=[\hat{{{\color[rgb]{0,0,0}\mathpzc{x}}}}_{1}^{\top}\cdots\hat{{{\color[rgb]{0,0,0}\mathpzc{x}}}}_{M}^{\top}]\in\mathbb{R}^{M\times n_{x}}$ and $Y^{M}=[\hat{{{\color[rgb]{0,0,0}\mathpzc{y}}}}_{1}\cdots\hat{{{\color[rgb]{0,0,0}\mathpzc{y}}}}_{M}]\in\mathbb{R}^{M}$ such that the Kullback-Leiber (KL) divergence between the posteriori distribution obtained from $\mathcal{D}^{M}$ and the original distribution based on $\mathcal{D}^{N}$ is minimal, while $M\ll N$.
@@
$\mathcal{L}(\mathcal{D}^{N},X^{M},\theta)=\frac{N}{2}\log(2\pi)+\frac{1}{2}Y^{\top}(W_{NN}+\sigma_{\epsilon}^{2}I_{N})Y^{N}\\ +\frac{1}{2}\log\det(W_{NN}+\sigma_{\epsilon}^{2}I_{N})+\frac{1}{2}\mathrm{tr}(K_{NN}-W_{NN})$ (8)
where $W_{NN}=K_{NM}K_{MM}^{-1}K_{MN}$ with $K_{MN}]_{i,j}=\kappa({\hat{{{\color[rgb]{0,0,0}\mathpzc{x}}}}_{i}},{{\color[rgb]{0,0,0}\mathpzc{x}}}_{j}),\;i\in\mathbb{I}_{1}^{M},\;j\in\mathbb{I}_{1}^{N}$, which is the covariance matrix between the pseudo inputs and all training inputs, $K_{MM}]_{i,j},\;i,j=\mathbb{I}_{1}^{M}$ is the covariance of the pseudo inputs. By maximizing (8), the hyperparameters of the GP and the pseudo inputs can be jointly obtained, i.e.
@@
After the training, the resulting posterior distribution at an arbitrary test point $\color[rgb]{0,0,0}\mathpzc{x}$$_{\star}$ is $\mathcal{N}(\mu_{\mathrm{p}}({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star}),\Sigma_{\mathrm{p}}({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star}))$, with
$\mu_{\mathrm{p}}({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star})$ $=K_{M}^{\top}({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star})Q_{MM}K_{MN}\sigma^{-2}_{\epsilon}Y^{N},$ (10a)
$\begin{split}\Sigma_{\mathrm{p}}({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star})&=\kappa({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star},{{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star})-W_{MM}({{\color[rgb]{0,0,0}\mathpzc{x}}}^{\star})+K_{M}^{\top}({{\color[rgb]{0,0,0}\mathpzc{x}}}^{\star})K_{MM}^{-1}K_{M}({{\color[rgb]{0,0,0}\mathpzc{x}}}^{\star}),\end{split}$ (10b)
where $K_{M}({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star})]_{i}=\kappa({{\color[rgb]{0,0,0}\mathpzc{x}}}_{i},{{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star}),\;i\in\mathbb{I}_{1}^{M}$ and $W_{MM}({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star})=K_{M}^{\top}({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star})K_{MM}^{-1}K_{M}({{\color[rgb]{0,0,0}\mathpzc{x}}}_{\star})$, which corresponds to a Nyström projection of the original GP to the pseudo inputs $X^{M}$. This formulation reduces the $\mathcal{O}(N^{3})$ computational complexity of the evaluation of the mean to $\mathcal{O}(M)$ and the training to $\mathcal{O}(NM^{2})$. In the following sections, SGPs will be utilized as the learning component of the proposed adaptive control approach.